OpenStudy (anonymous):

What is the slope of the line that passes through the points (-2,5) and (1,4)? A. -3 B. -2 C. -1/3 D.1/3

OpenStudy (astrophysics):

\[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} } \] here is the slope formula where your points are \[(x_1,y_1)~~~\text{and}~~~(x_2,y_2)\]

OpenStudy (anonymous):

slope equals \[\frac{ y^2-y^1 }{ x^2-x^1 }\]

OpenStudy (anonymous):

C. by inspection.

OpenStudy (anonymous):

Okay thank you

OpenStudy (anonymous):

Your welcome.

OpenStudy (blowery):

M=-1/3, or C in your multiple choice